# Fractional parts of powers of negative rationals

### 17 Mar 2026 v1

@@ 1 Introduction
Distribution modulo one of geometric progressions has been studied intensively. One intriguing open question considered by Mahler [Mahler] is the following: Does there exist a real number $\xi>0$ such that $\{\xi(3/2)^{n}\}<1/2$ for all $n\geq 0$? Here $\{x\}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}x-\lfloor x\rfloor$ denotes the fractional part of a real number $x$. Flatto, Lagarias, and Pollington [FLP, Theorem 1.4] proved that for any real number $\xi>0$ and integers $p>q\geq 1$ satisfying $\xi\notin\mathbb{Q}$ or $p/q\notin\mathbb{Z}$, we have
@@
It follows immediately from Theorem 1.1 that the image in $\mathbb{R}/\mathbb{Z}$ of the sequence $(\xi(-p/q)^{n})_{n\geq 0}$ is not contained in any interval of length less than $(1+r-r^{2})/p$. In subsequent work [LZ], we will examine the question whether the image can lie in an interval of length equal to $(1+r-r^{2})/p$. In particular, we will show that the bound in Theorem 1.1 is the best possible in the case $q=1$. The analogous question for powers of positive rational numbers was studied by several authors ([BD], [AFS], [Dubickas-small], [Dubickas-large]).
@@ Acknowledgments
We will call this formula (2.3) applied to the subsequences $(s_{n},\dots,s_{n+k-1})$ and $(s_{m},\dots,s_{m+k-1})$.
@@
The inequality (2.4) is a special case of [Dubickas, Lemma 3]. We include a proof for the sake of completeness. Assume that the set of limit values of $(S_{n}(r))_{n\geq 0}$ is contained in an interval $(A,B)$ of length <1<1. Then, up to shifting (and truncating) $S$, we may assume that $S_{n}(r)\in(A,B)$ for all $n\geq 0$. Let $\mu=\max_{n\geq 0}s_{n}$ and $\nu=\min_{n\geq 0}s_{n}$. Since $S$ is not ultimately periodic, we have $\mu>\nu$. By (2.3), $2>(1+r)(B-A)>\mu-\nu$. Thus $\mu-\nu=1$ and consequently $s_{n}\in\{\nu,\mu\}$ for all $n\geq 0$. We regard $S$ as a word $s_{0}s_{1}\dots$ over the alphabet $\{\nu,\mu\}$. Let $n\geq 0$. By Lemma 2.1, there exists a word $U=u_{1}\dots u_{n}$ of length $n$ such that $\mu U$ and $\nu U$ are subwords of $S$. By (2.3) applied to these subwords, $(1+r^{n+1})(B-A)>\mu-\nu=1$. Since $n\geq 0$ is arbitrary, we get $B-A\geq 1$. Contradiction. This finishes the proof of (2.4).
@@
Let $\mu=\max_{n\geq 0}s_{n}$ and $\nu=\min_{n\geq 0}s_{n}$. Since $S$ is not ultimately periodic, we have $\mu>\nu$. Moreover, there exists a subsequence $(\mu,\mu^{\prime})$ of $S$ with $\mu^{\prime}<\mu$. Similarly, there exists a subsequence $(\nu,\nu^{\prime})$ of $S$ with $\nu^{\prime}>\nu$. If $\mu-\nu\geq 2$, then, by (2.3) applied to the subsequences $(\mu,\mu^{\prime})$ and $(\nu,\nu^{\prime})$,
@@
Note that $S$ has a subword of the form $10a$. If $010$ is a subword of $S$, then, by (2.3) applied to these subwords, we have
@@
Let $k\geq 2$ be an even integer and let $W=w_{1}w_{2}\dots$ be a Sturmian word over the alphabet $\{0,1\}$. Let $z_{n}=2(k+w_{n})$ for $n\geq 1$ and let $z_{0}\geq 1$ be an integer. Consider the word $S=0^{z_{0}}110^{z_{1}}11\dots$. We will show that
@@
For $n\geq 0$, let $\sigma_{n}=-1+\sum_{i=0}^{n-1}(z_{n}+2)=-1+2kn+2\tau_{n}$, where $\tau_{n}=\sum_{i=0}^{n-1}w_{n}$. Then $s_{j}=1$ if and only if $s_{j}=\sigma_{n}$ or $s_{j}=\sigma_{n}-1$ for some $n\geq 1$. For $n\geq 1$, we have
(2.6) $S_{\sigma_{n}}(-r)=1-\sum_{i=n+1}^{\infty}r^{\sigma_{i}-\sigma_{n}-1}(1-r)=1-\sum_{i=n+1}^{\infty}r^{2(i-n)-1}(1-r)>1-\frac{r(1-r)}{1-r^{2}}=\frac{1}{1+r},$
@@
Here in (2.6) we used the estimate $\sigma_{i}-\sigma_{n}\geq 2k(i-n)>2(i-n)$ for $i\geq n+1$. By (2.6), $S_{\sigma_{n}-1}(-r)=1-rS_{\sigma_{n}}(-r)<S_{\sigma_{n}}(-r)$. Moreover, for $\sigma_{n-1}<j<\sigma_{n}$, $S_{j}(-r)=(-r)^{\sigma_{n}-j-1}S_{\sigma_{n}-1}(-r)$. Thus

### 24 Mar 2026 v2

@@ 1 Introduction
Distribution modulo one of geometric progressions has been studied extensively. One intriguing open question considered by Mahler [Mahler] is the following: Does there exist a real number $\xi>0$ such that $\{\xi(3/2)^{n}\}<1/2$ for all $n\geq 0$? Here $\{x\}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}x-\lfloor x\rfloor$ denotes the fractional part of a real number $x$. Flatto, Lagarias, and Pollington [FLP, Theorem 1.4] proved that for any real number $\xi>0$ and coprime integers $p>q>1$, we have
@@
It follows immediately from Theorem 1.1 that the image in $\mathbb{R}/\mathbb{Z}$ of the sequence $(\xi(-p/q)^{n})_{n\geq 0}$ is not contained in any interval of length less than $(1+r-r^{2})/p$. In subsequent work [LZ], we will examine the question whether the image can lie in an interval of length equal to $(1+r-r^{2})/p$. In particular, we will show that the bound in Theorem 1.1 is the best possible in the case $q=1$. The analogous question for powers of positive rational numbers was studied by several authors ([FLP], [Bugeaud-linear], [BD], [Dubickas-small], [Dubickas-large]).
@@ Acknowledgments
We will call this formula (2.3) applied to the finite subsequences $(s_{n},\dots,s_{n+k-1})$ and $(s_{m},\dots,s_{m+k-1})$.
@@
We include a proof of (2.4) for the sake of completeness. Assume that the set of limit values of $(S_{n}(r))_{n\geq 0}$ is contained in an interval $(A,B)$ of length <1<1. Then, up to shifting (and truncating) $S$, we may assume that $S_{n}(r)\in(A,B)$ for all $n\geq 0$. Let $\mu=\max_{n\geq 0}s_{n}$ and $\nu=\min_{n\geq 0}s_{n}$. Since $S$ is not ultimately periodic, we have $\mu>\nu$. By (2.3),
$2>(1+r)(B-A)>\mu-\nu.$
Thus $\mu-\nu=1$ and consequently $s_{n}\in\{\nu,\mu\}$ for all $n\geq 0$. We regard $S$ as a word $s_{0}s_{1}\dots$ over the alphabet $\{\nu,\mu\}$. Let $n\geq 0$. By Lemma 2.1, there exists a word $U=u_{1}\dots u_{n}$ of length $n$ such that $\mu U$ and $\nu U$ are subwords of $S$. By (2.3) applied to these subwords,
$(1+r^{n+1})(B-A)>\mu-\nu=1.$
Since $n\geq 0$ is arbitrary, we get $B-A\geq 1$. Contradiction. This finishes the proof of (2.4).
@@
Let $\mu=\max_{n\geq 0}s_{n}$ and $\nu=\min_{n\geq 0}s_{n}$. Since $S$ is not ultimately periodic, we have $\mu>\nu$. Moreover, there exists a finite subsequence $(\mu,\mu^{\prime})$ of $S$ with $\mu^{\prime}<\mu$. Similarly, there exists a finite subsequence $(\nu,\nu^{\prime})$ of $S$ with $\nu^{\prime}>\nu$. If $\mu-\nu\geq 2$, then, by (2.3) applied to the finite subsequences $(\mu,\mu^{\prime})$ and $(\nu,\nu^{\prime})$,
@@
Note that $S$ has a subword of the form $10a$, where $a\in\{0,1\}$. If $010$ is a subword of $S$, then, by (2.3) applied to these subwords, we have
@@
Let $k\geq 2$ be an even integer and let $W=w_{0}w_{1}\dots$ be a Sturmian word over the alphabet $\{0,1\}$. Let $z_{n}=k+2w_{n}$ for $n\geq 0$. Consider the word $S=0^{z_{0}}110^{z_{1}}11\dots$. We will show that
@@
For $n\geq 0$, let $\sigma_{n}=-1+\sum_{i=0}^{n-1}(z_{i}+2)=-1+(k+2)n+2\tau_{n}$, where $\tau_{n}=\sum_{i=0}^{n-1}w_{i}$. Then $s_{j}=1$ if and only if $s_{j}=\sigma_{n}$ or $s_{j}=\sigma_{n}-1$ for some $n\geq 1$. For $n\geq 1$, we have
(2.6) $S_{\sigma_{n}}(-r)=1-\sum_{i=n+1}^{\infty}r^{\sigma_{i}-\sigma_{n}-1}(1-r)>1-\sum_{i=n+1}^{\infty}r^{2(i-n)-1}(1-r)>1-\frac{r(1-r)}{1-r^{2}}=\frac{1}{1+r},$
@@
Here in (2.6) we used the estimate $\sigma_{i}-\sigma_{n}\geq(k+2)(i-n)>2(i-n)$ for $i\geq n+1$. We have $S_{\sigma_{n}-1}(-r)=1-rS_{\sigma_{n}}(-r)<S_{\sigma_{n}}(-r)$ by (2.6). Moreover, for $\sigma_{n-1}<j<\sigma_{n}$, $S_{j}(-r)=(-r)^{\sigma_{n}-j-1}S_{\sigma_{n}-1}(-r)$. Thus